The measure of each exterior angle of a regular polygon is 2/13 of the measure of each interior angles. Find the number of sides and diagonals of the polygon.
Answer by Alan3354(69443)

The measure of each exterior angle of a regular polygon is 2/13 of the measure of each interior angles. Find the number of sides and diagonals of the polygon.
--------------
I = Interior angle
E = exterior angle = 2I/13
-----
E + I = 180
I + 2I/13 = 180
15I/13 = 180
I/13 = 12
I = 156 degs
E = 180 - I = 24 degs
number of sides = 360/24 = 15